# Reflections to set-theoretic solutions of the Yang-Baxter equation 11footnotemark: 1

### v2

@@ 5.1 The reflection of the matched product of two solutions
Let us begin by recalling the essential notions contained in [2]. If $(S,r_{S})$ and $(T,r_{T})$ are two solutions, $\alpha:T\to\operatorname{Sym}\left(S\right)$ and $\beta:S\to\operatorname{Sym}\left(T\right)$ are maps, set $\alpha_{u}:=\alpha\left(u\right)$, for every $u\in T$, and $\beta_{a}:=\beta\left(a\right)$, for every $a\in S$. Then, the quadruple $\left(r_{S},r_{T},\alpha,\beta\right)$ is said to be a matched product system of solutions if the following conditions hold
@@
Let $(a,u),(b,v)\in S\times T$. We show the proof for the first components of the identities in $1$ and $2$, since the proof for the second ones will be similar, by exchanging the roles of $a$ and $u$, $b$ and $v$, $\mathcal{A}$ and $\mathcal{U}$, $\alpha$ and $\beta$, and $\kappa$ and $\omega$, and by using $\left(s6\right)$ instead of $\left(s5\right)$. The first component of $t\left(\left(a,u\right),\left(\kappa\left(b\right),\omega\left(v\right)\right)\right)$ is given by
@@
and the first component of $\left(\left(a,u\right),\left(b,v\right)\right)$ is
@@ 5.2 The reflection of the strong semilattice of solutions
Let us start by recalling the construction of strong semilattice of solutions provided in [3, Theorem 4.1]. Let $Y$ be a (lower) semilattice, $\left\{X_{\alpha}\ \left|\ \alpha\in Y\right.\right\}$ a family of disjoint sets indexed by $Y$, and $\left(X_{\alpha},r_{\alpha}\right)$ a solution, for every $\alpha\in Y$. For all $\alpha,\beta\in Y$ such that $\alpha\geq\beta$, let $\phi_{\alpha,\beta}:X_{\alpha}\to X_{\beta}$ be a map such that
@@ 5.3 The reflection of the twisted unions
In [15, Section 3.4], Etingof, Schedler, and Soloviev provided a technique to construct involutive non-degenerate solutions in the finite case, called the twisted union. A computer program generated all solutions, up to isomorphism, on sets of cardinality $|X|\leq 8$, and showed that almost all are twisted union of involutive non-degenerate solutions on smaller sets. We extend such a construction to arbitrary solutions.
@@
$g\rho_{\alpha(y)}=\rho_{y}g$
@@
Let $\left(X,u\right)$ be the solution associated to a left rack $\left(X,\triangleright\right)$ and $(Y,v)$ the solution associated with a right rack $\left(Y,\triangleleft\right)$. Moreover, let $\kappa\in\mathcal{K}(X,u)$ and $\omega\in\mathcal{K}(Y,v)$. Set $f=\alpha=\kappa$, $g=\beta=\omega$, and $Z=X\mathbin{\mathchoice{\ooalign{$\displaystyle\cup$\cr$\displaystyle\cdot$}}{\ooalign{$\textstyle\cup$\cr$\textstyle\cdot$}}{\ooalign{$\scriptstyle\cup$\cr$\scriptstyle\cdot$}}{\ooalign{$\scriptscriptstyle\cup$\cr$\scriptscriptstyle\cdot$}}}Y$, by 2.11 and 5.9, we obtain that the map $r:Z\times Z\to Z\times Z$ defined by
@@
$\omega(a)=\begin{cases}\kappa(a)&\text{if}\quad a\in X\\ \kappa^{\prime}(a)&\text{if}\quad a\in Y\end{cases}$
@@ A Reflections for solutions associated with skew braces
We gather numerical data computing all reflections for some bijective non-degenerate solutions. To this aim, we focus on the solutions associated with the algebraic structure of skew braces. For more details about these structures, we refer the reader to [19]. In particular, we implement the GAP software and devise an elementary algorithmic approach to compute all reflections for the solutions associated to the first skew braces in the database furnished by the YangBaxter package [29].
The algorithm proceeds as explained below.
Algorithm.
@@ 4. Output:
##\## Type $(B,+)$ $(B,\circ)$ Tot. $\lambda$-cen. $\rho$-inv. Both
@@ 6 $C_{6}$ $C_{6}$ 46656 0 0 46656
Table 1: $\lambda$-cen. and $\rho$-inv. reflections for solutions associated with skew braces until order $7$.
Type $(B,+)$ $(B,\circ)$ Tot. $\lambda$-cen. $\rho$-inv. Both
@@ 47 $C_{2}\times C_{2}\times C_{2}$ $D_{8}$ 1278 63 0 1
Table 2: $\lambda$-cen. and $\rho$-inv. reflections for solutions associated with skew braces of order $8$.
@@
##\## Type $(B,+)$ $(B,\circ)$ $\lambda$-inv. $\rho$-cen. All
@@ 6 $C_{6}$ $C_{6}$ 0 0 46656
Table 3: $\lambda$-inv. and $\rho$-cen. reflections for solutions associated with skew braces until order $7$.

### v3

@@ 5.1 The reflection of the matched product of two solutions
Let us begin by recalling the essential notions contained in [2]. If $(S,r_{S})$ and $(T,r_{T})$ are solutions that, hereinafter, we write as $r_{S}\left(a,b\right)=\left(\lambda_{a}\left(b\right),\rho_{b}\left(a\right)\right)$, for all $a,b\in S$, and $r_{T}\left(u,v\right)=\left(\lambda_{u}\left(v\right),\rho_{v}\left(u\right)\right)$, for all $u,v\in T$, and $\alpha:T\to\operatorname{Sym}\left(S\right)$ and $\beta:S\to\operatorname{Sym}\left(T\right)$ are maps, set $\alpha_{u}:=\alpha\left(u\right)$, for every $u\in T$, and $\beta_{a}:=\beta\left(a\right)$, for every $a\in S$. Then, the quadruple $\left(r_{S},r_{T},\alpha,\beta\right)$ is said to be a matched product system of solutions if the following conditions hold
@@
In the following lemmas, consistently with maps $t$ and $q$ introduced in Lemma 1.6, let us consider the maps $t,q$ from $\left(S\times T\right)\times\left(S\times T\right)$ defined by
$t\left(\left(a,u\right),\left(b,v\right)\right):=\lambda_{\lambda_{\left(a,u\right)}\left(b,v\right)}\left(\kappa\times\omega\right)\rho_{\left(b,v\right)}\left(a,u\right)$
and
$q\left(\left(a,u\right),\left(b,v\right)\right):=\rho_{\left(\kappa\times\omega\right)\rho_{\left(b,v\right)}\left(a,u\right)}\lambda_{\left(a,u\right)}\left(b,v\right),$
for all $\left(a,u\right),\left(b,v\right)\in S\times T$, where $\kappa:S\to S$ and $\omega:T\to T$ are maps.
@@
Let $(a,u),(b,v)\in S\times T$. We show the proof for the first components of the identities in $1$ and $2$, since the proof for the second ones will be similar, by exchanging the roles of $a$ and $u$, $b$ and $v$, $\mathcal{A}$ and $\mathcal{U}$, $\alpha$ and $\beta$, and $\kappa$ and $\omega$, and by using $\left(s6\right)$ instead of $\left(s5\right)$. We have that
$t\left(\left(a,u\right),\left(\kappa\left(b\right),\omega\left(v\right)\right)\right)$ $=\lambda_{\left(\alpha_{u}\lambda_{\overline{a}}\kappa(b),\beta_{a}\lambda_{\overline{u}}\omega(v)\right)}\kappa\times\omega\left(\alpha_{\overline{U}}\rho_{\alpha_{\overline{u}}\kappa(b)}(a),\beta^{-1}_{\overline{A}}\rho_{\beta_{a}\omega(v)}(u)\right)$
$=\lambda_{\left(\lambda_{a}\alpha_{\overline{u}}\kappa(b),\lambda_{u}\beta_{\overline{a}}\omega(v)\right)}\kappa\times\omega\left(\alpha_{\overline{U}}\rho_{\alpha_{\overline{u}}\kappa(b)}(a),\beta^{-1}_{\overline{A}}\rho_{\beta_{a}\omega(v)}(u)\right)$
$=\lambda_{\left(\mathcal{A},\,\mathcal{U}\right)}\left(\kappa\alpha_{\overline{U}}\rho_{\alpha_{\overline{u}}\kappa(b)}(a),\omega\beta^{-1}_{\overline{A}}\rho_{\beta_{a}\omega(v)}(u)\right)$
where in the second to last equality we use $\left(s5\right)$ and $\left(s6\right)$. Thus, the first component of $t\left(\left(a,u\right),\left(\kappa\left(b\right),\omega\left(v\right)\right)\right)$ is given by
@@
and the first component of $t\left(\left(a,u\right),\left(b,v\right)\right)$ is
@@ 5.2 The reflection of the strong semilattice of solutions
Let us start by recalling the construction of strong semilattice of solutions provided in [3, Theorem 4.1]. Let $Y$ be a (lower) semilattice, $\left\{X_{\alpha}\ \left|\ \alpha\in Y\right.\right\}$ a family of disjoint sets indexed by $Y$, and $\left(X_{\alpha},r_{\alpha}\right)$ a solution, for every $\alpha\in Y$. For all $\alpha,\beta\in Y$, we simply denote the product $\alpha\wedge\beta$ in $Y$ by the juxtaposition $\alpha\beta$. Moreover, for all $\alpha,\beta\in Y$ such that $\alpha\geq\beta$, let $\phi_{\alpha,\beta}:X_{\alpha}\to X_{\beta}$ be a map such that
@@ 5.3 The reflection of the twisted unions
In [16, Section 3.4], Etingof, Schedler, and Soloviev provided a technique to construct involutive non-degenerate solutions in the finite case, called the twisted union. A computer program generated all solutions, up to isomorphism, on sets of cardinality $|X|\leq 8$, and showed that almost all are twisted union of involutive non-degenerate solutions on smaller sets. We extend such a construction to arbitrary solutions.
@@
$g\lambda_{\beta\left(y\right)}=\lambda_{y}g$
@@
Let $\left(X,u\right)$ be the solution associated to a left rack $\left(X,\triangleright\right)$ and $(Y,v)$ the solution associated with a right rack $\left(Y,\triangleleft\right)$. Moreover, consider $\kappa\in C_{\operatorname{End}(X,\triangleright)}\left(\operatorname{LMlt}\left(X,\triangleright\right)\right)$ (hence $\kappa$ is a reflection for $(X,u)$ by 2.12) and $\omega=R_{y}$ a reflection for $(Y,v)$, with $y\in Y$. If we set $f=\alpha=\kappa$, $g=\beta=\omega$, and $Z=X\mathbin{\mathchoice{\ooalign{$\displaystyle\cup$\cr$\displaystyle\cdot$}}{\ooalign{$\textstyle\cup$\cr$\textstyle\cdot$}}{\ooalign{$\scriptstyle\cup$\cr$\scriptstyle\cdot$}}{\ooalign{$\scriptscriptstyle\cup$\cr$\scriptscriptstyle\cdot$}}}Y$ then, by 3.6 and 5.9, the map $r:Z\times Z\to Z\times Z$ defined by
@@
$\omega(a)=\begin{cases}\kappa(a)&\text{if}\quad a\in X\\ \kappa^{\prime}(a)&\text{if}\quad a\in Y\end{cases}$ (4)
@@
Below, we provide some classes of examples obtained by 5.15.
Example 5.16.
Let $\left(X,u\right)$ and $\left(Y,v\right)$ be equivalent solutions via a bijection $\varphi:X\to Y$. Let us denote by $\left(Z,r\right)$ the solution obtained through 5.9 by considering $\alpha=f$ and $\beta=g$ maps such that $\left(X,u\right)$ and $\left(Y,v\right)$ are equivalent via $\alpha$ and $\beta$ with themselves, respectively. Furthermore, let $\kappa\in\mathcal{K}\left(X,u\right)$ and consider $\kappa^{\prime}=\kappa_{\varphi}(=\varphi\kappa\varphi^{-1})$ that, by 1.2-$1$, is a reflection for $\left(Y,v\right)$. Then, applying 5.15, we deduce that the map $\omega:Z\to Z$ defined as in (4) is a reflection for $(Z,\,r)$.
Example 5.17.
Let us consider the twisted union $(Z,r)$ of $\left(X,u\right)$ and $\left(Y,v\right)$ in 5.10. According to 1.4, any maps $\kappa:X\to X$ and $\kappa^{\prime}:Y\to Y$ such that $\kappa(f\alpha)=(f\alpha)k$ and $\kappa^{\prime}(g\beta)=(g\beta)\kappa^{\prime}$ are reflections for $(X,u)$ and $(Y,v)$, respectively. Thus, by 5.15, the map $\omega:Z\to Z$ defined as in (4) is a reflection for $(Z,r).$
Example 5.18.
Let $(Z,r)$ be the twisted union of $(X,u)$ and $(Y,v)$ in 5.14, with $f=\alpha=\kappa\in\mathcal{K}(X,u)$ and $g=\beta=\kappa^{\prime}\in\mathcal{K}(Y,v)$. Then by 5.15 the map $\omega:Z\to Z$ defined as in (4) is easily a reflection for $(Z,\,r)$.
@@ A Reflections for solutions associated with skew braces
We gather numerical data computing all reflections for some bijective non-degenerate solutions. To this aim, we focus on the solutions associated with the well-studied algebraic structure of skew braces. For more details on these structures, we refer the reader to [20]. Reflections maps coming from skew braces merit further investigation which will be undertaken in future works. In particular, we implement the GAP software and devise an elementary algorithmic approach to compute all reflections for the solutions associated to the first skew braces in the database furnished by the YangBaxter package [30].
The algorithm proceeds as explained below:
@@ 4. Output:
##\## Type $(B,+)$ $(B,\circ)$ Tot. $\lambda$-centr. $\rho$-inv. Both
@@ 6 $C_{6}$ $C_{6}$ 46656 0 0 46656
Table 1: $\lambda$-centr. and $\rho$-inv. reflections for solutions associated with skew braces until order $7$.
Type $(B,+)$ $(B,\circ)$ Tot. $\lambda$-centr. $\rho$-inv. Both
@@ 47 $C_{2}\times C_{2}\times C_{2}$ $D_{8}$ 1278 63 0 1
Table 2: $\lambda$-centr. and $\rho$-inv. reflections for solutions associated with skew braces of order $8$.
@@
##\## Type $(B,+)$ $(B,\circ)$ $\lambda$-inv. $\rho$-centr. All
@@ 6 $C_{6}$ $C_{6}$ 0 0 46656
Table 3: $\lambda$-inv. and $\rho$-centr. reflections for solutions associated with skew braces until order $7$.
@@ Acknowledgments
We thank the referee for carefully reading our manuscript and the remarks and suggestions that helped us improve the paper.